# Learning Geospatial Region Embedding with Heterogeneous Graph

### v1

@@ F.5 Details of investigation of GeoHG with GNNExplainer
    The target region exhibits strong associations with societal entity nodes like "Food and Beverage" and "Shopping Mall", which are typically carbon-intensive due to factors such as energy usage (cooking, refrigeration), transportation (goods/personnel movement), and packaging consumption.
@@
    Notably, important hyperedges also connect the target region to relatively distant areas like #(40, 109) and #(47, 49). Despite geographical separation, these regions share similar social contexts, being linked to the "Food and Beverage" node, thereby providing relevant emission patterns to inform the prediction.
These interpretable insights validate GeoGraph’s effectiveness in capturing the intricate high-order dependencies between environmental conditions, urbanization factors, economic activities, and their synergistic impact on carbon footprints across regions. By encoding such high-order interactions through graph structures, GeoHG offers a powerful inductive bias tailored for modeling complex geospatial phenomena governed by higher-order environment-society couplings.
@@ Appendix G Qualitative Demonstration
In this section, we qualitatively evaluate the efficacy of our proposed approach. To begin with, we visually depict the regional dependency distribution by calculating the cosine similarity between embeddings of distinct regions, which are acquired through self-supervised training as outlined in Section 3.4. Moreover, we juxtapose the pertinent real-world information of regions exhibiting highly similar embeddings to scrutinize the capacity of our method to succinctly encapsulate the information and interrelationships within geospace.
@@
$S_{C}(E_{A},E_{B})=\cos(\theta)=\frac{\mathbf{E_{A}}\cdot\mathbf{E_{B}}}{\|\mathbf{E_{A}}\|\|\mathbf{E_{B}}\|}=\frac{\sum_{i=1}^{n}E_{Ai}E_{Bi}}{\sqrt{\sum_{i=1}^{n}E_{Ai}^{2}}\cdot\sqrt{\sum_{i=1}^{n}E_{Bi}^{2}}}$ (16)
@@
Refer to captionRefer to caption
Figure 12: Visualization of Region Similarity in Geospatial Embedding of Guangzhou. Darker regions indicate higher similarity.
@@
For Region (36,63), we find that it is an urban resident area with green land near a river, named Caohe Village. From the heatmap, we can see that our model connects other residential areas exhibiting similar characteristics in Guangzhou. For example, Shilou Village and Lyv Village share similar traits, with residential settlements bordering flowing rivers.
It is worth noting that, with the comprehensive representation of environmental, societal, and spatial information in the geospace and the specially designed graph structure, our models do not simplify the content region second-order region pair based solely on their intra-feature similarity. Instead, we consistently identify grouped high-order regions through the embedding results, rather than merely relying on individual high-similarity points.
@@ Appendix H More Discussion
Limitations and Future Directions. One limitation of our proposed method is that the complexity of the graph structure is directly proportional to the area size of the geospace, as the unit grid of the geospace is fixed to a $1km\times 1km$ square. This could present a challenge in worldwide large-scale applications, where the area of geospace may be extremely vast. In the future, we may explore the potential of modeling the geospace of the entire Earth through an adaptive grid partition and graph construction method. Furthermore, although we have emphasized and validated the efficacy and advantages of segmentation-based satellite imagery encoding, the upper limit of this approach is determined by the quality of segmentation results provided by third parties. Additionally, while we have demonstrated the effectiveness of our GeoHG in various real-world applications, further investigation is warranted to evaluate its performance in more diverse and complex scenarios.
Social Impacts. Efficient geospatial embedding holds considerable benefits for the broader geographic and spatial computing communities. By comprehensively representing intra-region features and inter-region correlations, our proposed GeoHG framework exhibits significant potential to effect meaningful change across various domains, especially within the realms of smart cities and geoscience. The interoperability and efficiency provided by GeoHG facilitate a deeper understanding of complex geospace and the underlying mixed-order correlations inherent in the space. This comprehensive representation of regional spatial regions empowers stakeholders to monitor cities and environments more effectively, thereby making informed decisions that ultimately enhance individual and communal quality of life while fostering more resilient and sustainable environments.
Moreover, the remarkable data efficiency of GeoHG enables the community to investigate regional fine-grained climates with limited resources. For instance, in Section 4.3, we showcase its exceptional performance in predicting fine-grained ($1km\times 1km$) indicators for a large area ($16,401km^{2}$) using only 863 monitoring points. As human society progresses and the global environment changes, regional extreme climates, such as urban heat islands and local air pollution, continue to pose significant economic, environmental, and health challenges. Thus, effective fine-grained regional climate monitoring becomes increasingly paramount. It is noticed that the millions of lives lost annually due to local extreme heat and air pollution, we believe the enhanced geospatial region representation and data efficiency facilitated by our approach will support more effective human health protection measures and inform geopolitical decision-making, promoting improved urban and environmental management as well as extreme climate mitigation strategies.

### v2

@@ F.2 GNNExplainer
    The target region exhibits strong associations with societal entity nodes like ”Food and Beverage” and ”Shopping Mall”, which are typically carbon-intensive due to factors such as energy usage (cooking, refrigeration), transportation (goods/personnel movement), and packaging consumption.
@@
    Notably, important hyperedges also connect the target region to relatively distant areas like #(40, 109) and #(47, 49). Despite geographical separation, these regions share similar social contexts, being linked to the ”Food and Beverage” node, thereby providing relevant emission patterns to inform the prediction.
These interpretable insights validate GeoHG’s effectiveness in capturing the intricate high-order dependencies between environmental conditions, urbanization factors, economic activities, and their synergistic impact on carbon footprints across regions. By encoding such high-order interactions through graph structures, GeoHG offers a powerful inductive bias tailored for modeling complex geospatial phenomena governed by higher-order environment-society couplings.
@@ Appendix G Qualitative Demonstration
Figure 12: Visualization of Regions’ Similarity in Learned Embedding of Guangzhou. Darker regions indicate higher similarity with the region of interest.
@@
$S_{C}(E_{A},E_{B})=\frac{\mathbf{E_{A}}\cdot\mathbf{E_{B}}}{\|\mathbf{E_{A}}\|\|\mathbf{E_{B}}\|}=\frac{\sum_{i=1}^{n}E_{Ai}E_{Bi}}{\sqrt{\sum_{i=1}^{n}E_{Ai}^{2}}\cdot\sqrt{\sum_{i=1}^{n}E_{Bi}^{2}}}$ (16)
@@
For Region (36,63), we find that it is an urban resident area with green land near a river, named Caohe Village. From the heatmap, we can see that our model connects other residential areas exhibiting similar characteristics in Guangzhou. For example, Shilou Village and Lyv Village share similar traits, with residential settlements bordering flowing rivers. It is worth noting that, with the comprehensive representation of environmental, societal, and spatial information in the geospace and the specially designed graph structure, our models do not simplify the content region second-order region pair based solely on their intra-feature similarity. Instead, we consistently identify grouped high-order regions through the embedding results, rather than merely relying on individual high-similarity points.
@@ Appendix H More Discussion
H.1 Limitations and Future Directions.
One limitation of our proposed method is that the complexity of the graph structure is directly proportional to the area size of the geospace, as the unit grid of the geospace is fixed to a $1km\times 1km$ square. This could present a challenge in worldwide large-scale applications, where the area of geospace may be extremely vast. In the future, we may explore the potential of modeling the geospace of the entire Earth through an adaptive grid partition and graph construction method. Furthermore, although we have emphasized and validated the efficacy and advantages of segmentation-based satellite imagery encoding, the upper limit of this approach is determined by the quality of segmentation results provided by third parties. Additionally, while we have demonstrated the effectiveness of our GeoHG in various real-world applications, further investigation is warranted to evaluate its performance in more diverse and complex scenarios.
H.2 Social Impacts.
Efficient socioeconomic indicator inference holds considerable benefits for the broader geographic and spatial computing communities. By comprehensively representing intra-region features and inter-region correlations, our proposed GeoHG framework exhibits significant potential to effect meaningful change across various domains, especially within the realms of smart cities and geoscience. The interoperability and efficiency provided by GeoHG facilitate a deeper understanding of complex geospace and the underlying mixed-order correlations inherent in the space. This comprehensive representation of regional spatial regions empowers stakeholders to monitor cities and environments more effectively, thereby making informed decisions that ultimately enhance individual and communal quality of life while fostering more resilient and sustainable environments.
Moreover, the remarkable data efficiency of GeoHG enables the community to investigate regional fine-grained climates with limited resources. For instance, in our experiment, we showcase its exceptional performance in predicting fine-grained ($1km\times 1km$) indicators for a large area ($16,401km^{2}$) using only 863 monitoring points. As human society progresses and the global environment changes, regional extreme climates, such as urban heat islands and local air pollution, continue to pose significant economic, environmental, and health challenges. Thus, effective fine-grained regional climate monitoring becomes increasingly paramount. It is noticed that the millions of lives lost annually due to local extreme heat and air pollution, we believe the enhanced geospatial region representation and data efficiency facilitated by our approach will support more effective human health protection measures and inform geopolitical decision-making, promoting improved urban and environmental management as well as extreme climate mitigation strategies.